An athlete runs 100m in 12s. What is his speed in km/hr?
1
Expert's answer
2019-04-19T10:27:53-0400

"S=100m=\frac{100}{1000}=0.1km""t=12s=\frac{12}{3600} h"

"v=\frac{S}{t}=\frac{0.1\times3600}{12}=30 \frac{km}{h}"